What is the wavelength of the photon?
asot Edge 88049282 bers into Equations t previous | 6 of 6 I return to Part C The equation for photon energy, E, is uations. E- he and/or where h 6626-10*J s(Plancks constant) and c 2.99 10 m/s (the speed of light) What is the wavelength, λ of a photon that has an energy of E-368×10-10 J ? Express your answer numerically in meters hat can as area, A Hints a variable allows uS as we Submit My Answers Give Up y its value ituted for its Incorrect, One attempt remaining: Try Again Units in Parts A through C, the quantities were already given with the appropriate units However, you should always determine which unit is most appropriate for each quantity, then do any necessary conversions before substituting numbers into the formula Part D gives an example of this type of situation
Answer #1

E+hc/λ

λ=hc/E

Plug values,

λ =(6.626*10^-34*2.99*10^8)/(3.68*10^-19)

λ=5.38*10^-7 m